OpenStudy (anonymous):

find the average velocity in (m/s) of a bicycler that start 150 meters north of town and is 1200 meters north of town after 30.0 minuets.

OpenStudy (anonymous):

Speed= distance/Time Speed= (1200-150)/30 Speed= 1185Meters/30 min Speed= 39.5 Meters/1min Speed= 0.65833333/ per second

OpenStudy (adamk):

Since when is 1200 - 150 = 1185...? It should be: Velocity = (1200-150)/30 Velocity = 1050m/30min Velocity = 35m/1min Velocity = 0.58m/s